# Information-Theoretic Opacity-Enforcement in Markov Decision Processes

### v1

@@ 1 Introduction
$\underbrace{\text{initial uncertainty}}_{H(Z)-\text{``entropy before observation''}}-\underbrace{\text{remaining uncertainty}}_{H(Z|Y)-\text{`` entropy after observation''}}$
@@
Related work Opacity was originally introduced by Mazaré (2004), referring to concealing secrets from an observer. This work has led to various qualitative analyses of opacity within discrete event systems (DESs), where the system dynamics are governed by both controllable and uncontrollable events, and its information leakage is also defined by observable and non-observable events. In qualitative analysis, a trajectory is opaque if it satisfies some secret property and is indistinguishable from at least one trajectory that does not satisfy the secret. A system is opaque if it generates only an opaque trajectory.
Different types of qualitative opacity have been studied, including state-based, which requires the secret behavior of the system (i.e., the membership of its initial/current/past state to the set) to remain opaque Saboori and Hadjicostis (2007, 2013a); Han et al. (2023); Saboori and Hadjicostis (2013b, 2009); Yin et al. (2019); language-based Dubreil et al. (2008); Lin (2011); Shi et al. (2023), which aims to hide a set of secret executions; and “model-basedKeroglou and Hadjicostis (2016), aiming to prevent the observer from finding out the true model of the system among several candidates.
While qualitative opacity evaluates whether a system is opaque, a quantitative analysis of opacity is essential for measuring the degree of opacity in a system. Bérard et al. (2015a) defines quantitative opacity as the probability of generating an opaque run in a stochastic system. A related notion to quantitative opacity is covertness Marzouqi and Jarvis (2004, 2005), which measures the distance to the risk of being observed during the planning of a path to a predefined destination. This can be viewed as a model-based opacity, as it makes the observer uncertain if a covert agent is present or not. Recent work Ma et al. (2023) studies how covert planning can leverage the coupling of stochastic dynamics and observational noises to make the covert agent’s policy indistinguishable from a nominal policy.
An alternative approach involves the utilization of the value of information (VOI). Khouzani and Malacaria (2017) employed entropy as a measure of VOI to enforce opacity in channel design problems. The channel design problem can be viewed as a one-step decision-making problem. Recent work adds a linear temporal logic constraint to the entropy maximization problem on MDP Savas et al. (2020). Another recent work Molloy and Nair (2023) developed a method for obfuscating state trajectories in partially observable Markov decision processes (POMDPs) and shows that the conditional entropy of the state trajectory can be expressed into a cumulative sum of entropy terms, permitting the use of POMDP solvers to maximize/minimize the conditional entropy.
@@
Our contribution Our work can be seen as a generalization of minimal information leakage channel design from one-stage decision-making to sequential decision-making in Markov decision processes, concerning state-based opacity. We establish the last-state opacity and initial-state opacity by leveraging Shannon entropy Shannon (1949) as a symmetric measure of opacity. Different from Molloy and Nair (2023), we do not assume that the observer has access to control inputs. Thus, the conditional entropy of state-based opacity properties cannot be constructed as a cumulative sum of entropy terms. We propose a novel approach to compute/estimate the gradient of the conditional entropy with respect to (w.r.t.) the parameters of the control policy, leveraging the forward-backward algorithm in HMMs. Building on the gradient estimation/computation, we can then formulate the optimal opacity enforcement planning problem, subject to task constraints, into a constrained optimization and employ primal-dual gradient-based optimization to compute a (locally)-optimal policy that maximizes the opacity while satisfying a constraint on the total return. The correctness of the proposed methods is experimentally validated.
@@ 2.1 Preliminaries
Notation
The set of real numbers is denoted by $\mathbf{R}$. Random variables will be denoted by capital letters, and their realizations by lowercase letters (e.g., $X$ and $x$). The probability mass function (pmf) of a discrete random variable $X$ will be written as $p(x)$, the joint pmf of $X$ and $Y$ as $p(x,y)$, and the conditional pmf of $X$ given $Y=y$ as $p(x|y)$ or $p(x|Y=y)$. The sequence of random variables and their realizations with length $T$ are denoted as $X_{[0:T]}$ and $x_{[0:T]}$, respectively. Given a finite and discrete set $\mathcal{S}$, let $\mathcal{D}(\mathcal{S})$ be the set of all probability distributions over $\mathcal{S}$. The set $\mathcal{S}^{T}$ denotes the set of sequences with length $T$ composed of elements from $\mathcal{S}$ and $\mathcal{S}^{\ast}$ denotes the set of all finite sequences generated from $\mathcal{S}$.
@@
The problem is modeled as a MDP $M=\langle\mathcal{S},\mathcal{A},P,\mu_{0},R,\gamma\rangle$ where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is a finite set of actions, $P:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}(\mathcal{S})$ is a probabilistic transition function and $P(s^{\prime}|s,a)$ is the probability of reaching state $s^{\prime}$ given that action $a$ is taken at the state $s$. $\mu_{0}$ is the initial state distribution. The planning objective is described by a reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbf{R}$. $\gamma\in[0,1]$ is the discount factor.
@@
Note that the assumption on non-observable actions can be easily relaxed, by augmenting the state space with $\mathcal{S}\cup\mathcal{S}\times\mathcal{A}$ and defining the state-observation function over the augmented state space. This assumption is made for clarity.
For an MDP $M$ and P2’s observation function $\mathsf{Obs}$, a policy $\pi$ induces a discrete stochastic process $\{S_{t},O_{t},t\in\mathbf{N}\}$ where each $S_{t}\in\mathcal{S}$ and $O_{t}\in\mathcal{O}$. In the case of a Markov policy,
@@
Next, we will introduce some basic definitions in information theory that are used to quantify the opacity of a policy concerning a secret.
@@ 3.1 Primal-Dual Policy Gradient for Constrained Minimal Information Leakage
$\max_{\theta}\min_{\lambda\geq 0}L(\theta,\lambda)=H(Z_{T}|Y;\theta)+\lambda(V(\mu_{0},\theta)-\delta)$ (3)
@@ 3.2 Computing the Gradient of Conditional Entropy
where the probability measure $P_{\theta}\coloneqq P^{\mathcal{M}_{\pi_{\theta}}}$. The discrete conditional entropy of a binary random variable has a property that $0\leq H(Z_{T}|Y;\theta)\leq 1$. Taking the gradient of $H(Z_{T}|Y;\theta)$ w.r.t. the policy parameter $\theta$. By using a trick that $\nabla_{\theta}P(y)=P_{\theta}(y)\nabla_{\theta}\log P(y)$ and the property of conditional probability, we have
@@
$+P_{\theta}(z_{T},y)\nabla_{\theta}\log_{2}P_{\theta}(z_{T}|y)\Big{]}.$
@@
$P_{\theta}(z_{T}|y)\nabla_{\theta}P_{\theta}(y)\log_{2}P_{\theta}(z_{T}|y)+P_{\theta}(y)\frac{\nabla_{\theta}P_{\theta}(z_{T}|y)}{\log 2}\Big{]}.$
@@
$+P_{\theta}(z_{T}|y)\log_{2}P_{\theta}(z_{T}|y)\nabla_{\theta}\log P_{\theta}(y)+\frac{\nabla_{\theta}P_{\theta}(z_{T}|y)}{\log 2}\Big{]}$
@@
which is the joint probability of receiving observation $o_{[0:t]}$ and arriving at state $j$ at the $t$-th time step. By definition of $\alpha$ messages and the law of total probability, the probability of receiving observation $y$ is $P_{\theta}(y)=\sum_{i\in\mathcal{S}}\alpha(i,\theta)$.
@@
After computing $\nabla_{\theta}P_{\theta}(Z_{T}=1|y),\nabla_{\theta}P_{\theta}(Z_{T}=0|y)$, and $\nabla_{\theta}P(y)$, we can obtain the value of $\nabla_{\theta}H(Z_{T}|Y;\theta)$ by equation (LABEL:eq:HMM_gradient_entropy). It is noted that though $\mathcal{O}^{T}$ is a finite set of observations, it is combinatorial and may be too large to enumerate. To mitigate this issue, we can employ sample approximations to estimate $\nabla_{\theta}H_{\theta}(Z_{T}|Y)$: Given $M$ sequences of observations $\{y_{1},\dots,y_{M}\}$, we can approximate $H(Z_{T}|Y;\theta)$ by
$H(Z_{T}|Y;\theta)\approx-\frac{1}{M}\sum_{k=1}^{M}\sum_{z_{T}\in\{0,1\}}P_{\theta}(z_{T}|y_{k})\log_{2}P_{\theta}(z_{T}|y_{k}).$ (15)
@@
$+P_{\theta}(z_{T}|y_{k})\log_{2}P_{\theta}(z_{T}|y_{k})\nabla_{\theta}\log P_{\theta}(y_{k})+\frac{\nabla_{\theta}P_{\theta}(z_{T}|y_{k})}{\log 2}\big{]}$
@@ 4 Synthesizing Maximally Opacity-Enforcement Controllers For Initial-state Opacity
$\max_{\theta}\min_{\lambda\geq 0}L(\theta,\lambda)=H(S_{0}|Y;\theta)+\lambda(V(s_{0},\theta)-\delta)$ (19)
@@
The computations of $P_{\theta}(\theta)$ and $\nabla_{\theta}\log P_{\theta}(y)$ are the same as these for the last-state opacity (see section 3.2). In addition, for the initial-state opacity, we need to obtain the $P_{\theta}(s_{0}|y)$ and its derivative w.r.t. $\theta$.
@@
$\nabla_{\theta}P_{\theta}(y|i)=\nabla_{\theta}\beta_{0}(i,\theta)=\\ \sum_{j\in\mathcal{S}}[P_{\theta}(i,j)b_{j}(o_{t+1})\nabla_{\theta}\beta_{t+1}(j,\theta)\\ +b_{j}(o_{t+1})\beta_{t+1}(j,\theta)\nabla_{\theta}P_{\theta}(i,j)],$ (25)
The computation of $\nabla_{\theta}P_{\theta}(i,j)$ is in equation (14). In this way, $H(S_{0}|Y;\theta)$ and $\nabla_{\theta}H(S_{0}|Y;\theta)$ can be computed exactly. Again, since $y\in\mathcal{O}^{T}$ is combinatorially many, we employ sample approximation to estimate the value. Given $M$ trajectories $\{y_{1},\dots,y_{M}\}$, the entropy term $H(S_{0}|Y;\theta)$ can be approximated by
$H(S_{0}|Y;\theta)\approx-\frac{1}{M}\sum_{k=1}^{M}\sum_{s_{0}\in\mathcal{S}}P_{\theta}(s_{0}|y_{k})\log_{2}P_{\theta}(s_{0}|y_{k}).$
And, the gradient term $\nabla_{\theta}H(S_{0}|Y;\theta)$ can be approximated by
@@ 5 Experiment Evaluation
The question marks on the grid represent the secret states for P1, while the flags denote the goal states for P1. We set the reward of reaching a goal to be 0.1 and the constraint that the total return is greater than or equals $\delta=0.3$, and the horizon $T=10$. Note that in this finite horizon, the robot can repeatedly visit the goal state.
@@
$H(s;\theta)=-\frac{1}{1-\gamma}E_{s\sim d_{\pi_{\theta}}}\Big{[}\sum_{a\in\mathcal{A}}\pi_{\theta}(a|s)\log\pi_{\theta}(a|s)\Big{]},$
@@
We compare our method with entropy-regularized MDPs solved with different $\tau$ values, selecting $10$ values from $\tau=0.01$ to $\tau=0.1$. The following graphs (Figure 3) illustrate the results. Note that when $\tau=0.1$, the policy approaches a random policy, making it impractical to further increase opacity by raising $\tau$.
The results in Figure 3 indicate that maximizing last-state opacity and initial-state opacity cannot be achieved using standard entropy-regularized MDP. In Figure 3(a), we observed that despite the increasing value for $\tau$, the conditional entropy $H(Z_{T}|Y)$ under the baseline algorithm remains below 0.4 and below the value obtained with the proposed method, indicating that the observer has more certainty in predicting if a secret state is visited. The value of the policy decreases as $\tau$ increases, and when $\tau\geq 0.03$, the entropy-regulated policy fails to satisfy the constraint on the total return. Our method outperforms this baseline by attaining the value of $0.3$ with a last-state opacity of $0.930$. Figure

### v2

@@ 1 Introduction
$\underbrace{\text{initial uncertainty}}_{H(Z)-\text{``entropy before observation''}}-\underbrace{\text{remaining uncertainty}}_{H(Z|Y)-\text{`` entropy after observation''}}.$
@@
Related Work Opacity was originally introduced by Mazaré Mazaré (2004), referring to concealing secrets from an observer. This work has led to qualitative opacity analyses within discrete event systems (DESs), where the system dynamics are governed by both controllable and uncontrollable events, and its information leakage is also defined by observable and non-observable events. In qualitative analysis, a trajectory is opaque if it satisfies some secret property and is indistinguishable from at least one trajectory that does not satisfy the secret. Furthermore, a system is opaque if it generates only an opaque trajectory. Different types of qualitative opacity have been studied, including state-based, which requires the secret behavior of the system (i.e., the membership of its initial/current/past state to the set) to remain opaque Saboori and Hadjicostis (2007, 2013a); Han et al. (2023); Saboori and Hadjicostis (2013b, 2009); Yin et al. (2019); language-based Dubreil et al. (2008); Lin (2011); Shi et al. (2023), which aims to hide a set of secret executions; and “model-basedKeroglou and Hadjicostis (2016), aiming to prevent the observer from finding out the true model of the system among several candidates. In multi-agent deterministic systems, cooperative privacy-preserving planning has been investigated Maliah et al. (2017); Shani (2018) to achieve a global coordinated plan without revealing local private information.
A quantitative analysis of opacity is essential for measuring the degree of opacity in a system. Bérard et al. (2015a) defines quantitative opacity as the probability of generating an opaque run in a stochastic system. A related notion to quantitative opacity is covertness Marzouqi and Jarvis (2004, 2005), which can be viewed as a model-based opacity, as it makes the observer uncertain if a covert agent is present or not. Recent work Ma et al. (2024) developed covert planner that leverages the coupling of stochastic dynamics and observational noises to make the covert agent’s policy indistinguishable from a nominal policy. Information-theoretic opacity has been studied in channel design problems, which can be viewed as a one-step decision-making problemKhouzani and Malacaria (2017). Another recent work Molloy and Nair (2023) developed a method for obfuscating state trajectories in partially observable Markov decision processes (POMDPs) and shows that the causal conditional entropy of the state trajectory can be expressed into a cumulative sum, permitting the use of POMDP solvers. Our work is also related to observer-aware MDP Miura and Zilberstein (2021) in the sense that one of the reward functions depends on the observer’s belief (of the secret). However, our solution does not employ POMDP solvers.
@@
Our Contribution Our work can be seen as a generalization of minimal information leakage channel design from one-stage decision-making to sequential decision-making in Markov decision processes, concerning state-based opacity. We establish the last-state opacity and initial-state opacity by leveraging Shannon entropy Shannon (1949) as a symmetric measure of opacity. Different from Molloy and Nair Molloy and Nair (2023), we do not assume that the observer has access to control inputs. Thus, the conditional entropy of state-based opacity properties cannot be constructed as a cumulative sum of entropy terms. We propose a novel approach to compute/estimate the gradient of the conditional entropy with respect to (w.r.t.) the parameters of the control policy, leveraging the forward-backward algorithm in HMMs. Building on the gradient estimation/computation, we can then formulate the optimal opacity enforcement planning problem, subject to task constraints, into a constrained optimization and employ primal-dual gradient-based optimization to compute a (locally)-optimal policy that maximizes the opacity while satisfying a constraint on the total return. The correctness of the proposed methods is experimentally validated.
@@ 2.1 Preliminaries
Notation The set of real numbers is denoted by $\mathbf{R}$. Random variables will be denoted by capital letters, and their realizations by lowercase letters (e.g., $X$ and $x$). The probability mass function (pmf) of a discrete random variable $X$ will be written as $p(x)$, the joint pmf of $X$ and $Y$ as $p(x,y)$, and the conditional pmf of $X$ given $Y=y$ as $p(x|y)$ or $p(x|Y=y)$. The sequence of random variables and their realizations with length $T$ are denoted as $X_{[0:T]}$ and $x_{[0:T]}$, respectively. Given a finite and discrete set $\mathcal{S}$, let $\mathcal{D}(\mathcal{S})$ be the set of all probability distributions over $\mathcal{S}$. The set $\mathcal{S}^{T}$ denotes the set of sequences with length $T$ composed of elements from $\mathcal{S}$, and $\mathcal{S}^{\ast}$ denotes the set of all finite sequences generated from $\mathcal{S}$.
@@
The problem is modeled as an MDP $M=\langle\mathcal{S},\mathcal{A},P,\mu_{0},R,\gamma\rangle$ where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is a finite set of actions, $P:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{D}(\mathcal{S})$ is a probabilistic transition function and $P(s^{\prime}|s,a)$ is the probability of reaching state $s^{\prime}$ given that action $a$ is taken at the state $s$, and $\mu_{0}$ is the initial state distribution. The planning objective is described by a reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbf{R}$. $\gamma\in[0,1]$ is the discount factor.
@@
Note that the assumption on non-observable actions can be easily relaxed, by augmenting the state space with $\mathcal{S}\cup(\mathcal{S}\times\mathcal{A})$ and defining the state-observation function over the augmented state space. This assumption is made for clarity.
For an MDP $M$ and P2’s observation function $\mathsf{Obs}$, a policy $\pi$ induces a discrete stochastic process $\{S_{t},O_{t},t\in\mathbf{N}\}$ where each $S_{t}\in\mathcal{S}$ and $O_{t}\in\mathcal{O}$. For a Markov policy,
@@
Next, we will introduce some basic definitions in information theory that are used to quantify the opacity of a policy.
@@ 3.1 Primal-Dual Policy Gradient for Constrained Minimal Information Leakage
$\max_{\theta}\min_{\lambda\geq 0}L(\theta,\lambda)=H(Z_{T}|Y;\theta)+\lambda(V(\mu_{0},\theta)-\delta),$ (3)
@@ 3.2 Computing the Gradient of Conditional Entropy
where the probability measure $P_{\theta}\coloneqq P^{\mathcal{M}_{\pi_{\theta}}}$. The discrete conditional entropy of a binary random variable has a property that $0\leq H(Z_{T}|Y;\theta)\leq 1$. We take the gradient of $H(Z_{T}|Y;\theta)$ w.r.t. the policy parameter $\theta$. By using a trick that $\nabla_{\theta}P(y)=P_{\theta}(y)\nabla_{\theta}\log P_{\theta}(y)$ and the property of conditional probability, we have
@@
$+P_{\theta}(z_{T},y)\nabla_{\theta}\log_{2}P_{\theta}(z_{T}|y)\Big{]}$
@@
$P_{\theta}(z_{T}|y)\nabla_{\theta}P_{\theta}(y)\log_{2}P_{\theta}(z_{T}|y)+P_{\theta}(y)\frac{\nabla_{\theta}P_{\theta}(z_{T}|y)}{\log 2}\Big{]}$
@@
$+P_{\theta}(z_{T}|y)\log_{2}P_{\theta}(z_{T}|y)\nabla_{\theta}\log P_{\theta}(y)+\frac{\nabla_{\theta}P_{\theta}(z_{T}|y)}{\log 2}\Big{]}.$
@@
which is the joint probability of receiving observation $o_{[0:t]}$ and arriving at state $j$ at the $t$-th time step. By definition of $\alpha$ messages and the law of total probability, the probability of receiving observation $y$ is $P_{\theta}(y)=\sum_{i\in\mathcal{S}}\alpha_{T}(i,\theta)$.
@@
After computing $\nabla_{\theta}P_{\theta}(Z_{T}=1|y),\nabla_{\theta}P_{\theta}(Z_{T}=0|y)$, and $\nabla_{\theta}P(y)$, we can obtain the value of $\nabla_{\theta}H(Z_{T}|Y;\theta)$ by equation (LABEL:eq:HMM_gradient_entropy). It is noted that, though $\mathcal{O}^{T}$ is a finite set of observations, it is combinatorial and may be too large to enumerate. To mitigate this issue, we can employ sample approximations to estimate $\nabla_{\theta}H_{\theta}(Z_{T}|Y)$: Given $M$ sequences of observations $\{y_{1},\dots,y_{M}\}$, we can approximate $H(Z_{T}|Y;\theta)$ by
$H(Z_{T}|Y;\theta)\approx-\frac{1}{M}\sum_{k=1}^{M}\sum_{z_{T}\in\{0,1\}}P_{\theta}(z_{T}|y_{k})\log_{2}P_{\theta}(z_{T}|y_{k}),$ (15)
@@
$+P_{\theta}(z_{T}|y_{k})\log_{2}P_{\theta}(z_{T}|y_{k})\nabla_{\theta}\log P_{\theta}(y_{k})+\frac{\nabla_{\theta}P_{\theta}(z_{T}|y_{k})}{\log 2}\big{]}.$
@@ 4 Synthesizing Maximally Opacity-Enforcement Controllers For Initial-state Opacity
$\max_{\theta}\min_{\lambda\geq 0}L(\theta,\lambda)=H(S_{0}|Y;\theta)+\lambda(V(s_{0},\theta)-\delta),$ (19)
@@
The computations of $P_{\theta}(\theta)$ and $\nabla_{\theta}\log P_{\theta}(y)$ are the same as those for the last-state opacity (see Section 3.2). In addition, for the initial-state opacity, we need to obtain the $P_{\theta}(s_{0}|y)$ and its derivative w.r.t. $\theta$.
@@
$\nabla_{\theta}P_{\theta}(y|i)=\nabla_{\theta}\beta_{0}(i,\theta)=\\ \sum_{j\in\mathcal{S}}[P_{\theta}(i,j)b_{j}(o_{t+1})\nabla_{\theta}\beta_{t+1}(j,\theta)\\ +b_{j}(o_{t+1})\beta_{t+1}(j,\theta)\nabla_{\theta}P_{\theta}(i,j)].$ (25)
The computation of $\nabla_{\theta}P_{\theta}(i,j)$ is in equation (14). In this way, $H(S_{0}|Y;\theta)$ and $\nabla_{\theta}H(S_{0}|Y;\theta)$ can be computed exactly. Again, since $y\in\mathcal{O}^{T}$ is combinatorially many, we employ sample approximation to estimate the value. Given $M$ trajectories $\{y_{1},\dots,y_{M}\}$, the entropy term $H(S_{0}|Y;\theta)$ and the gradient term $\nabla_{\theta}H(S_{0}|Y;\theta)$ can be approximated by
$H(S_{0}|Y;\theta)\approx-\frac{1}{M}\sum_{k=1}^{M}\sum_{s_{0}\in\mathcal{S}}P_{\theta}(s_{0}|y_{k})\log_{2}P_{\theta}(s_{0}|y_{k}),$
and
@@ 5 Experiment Evaluation
The question marks on the grid represent the secret states for P1, while the flags denote the goal states for P1. We set the reward of reaching a goal to be 0.1 and the constraint that the total return is greater than or equals $\delta=0.3$, and the horizon $T=10$. Note that, in this finite horizon, the robot can repeatedly visit the goal state.
@@
$H(s;\theta)=-\frac{1}{1-\gamma}E_{s\sim d_{\pi_{\theta}}}$
$\Big{[}\sum_{a\in\mathcal{A}}\pi_{\theta}(a|s)\log\pi_{\theta}(a|s)\Big{]},$
@@
We compare our method with entropy-regularized MDPs solved with different $\tau$ values, selecting $10$ values from $\tau=0.01$ to $\tau=0.1$. The following graphs (Figure 3) illustrate the results. Note that, when $\tau=0.1$, the policy approaches a random policy, making it impractical to further increase opacity by raising $\tau$.
The results in Figure 3 indicate that maximizing last-state opacity and initial-state opacity cannot be achieved using standard entropy-regularized MDP. In Figure 3(a), we observed that, despite the increasing value for $\tau$, the conditional entropy $H(Z_{T}|Y)$ under the baseline algorithm remains below 0.4 and below the value obtained with the proposed method, indicating that the observer has more certainty in predicting if a secret state is visited. The value of the policy decreases as $\tau$ increases, and when $\tau\geq 0.03$, the entropy-regulated policy fails to satisfy the constraint on the total return. Our method outperforms this baseline by attaining the value of $0.3$ with a last-state opacity of $0.930$. Figure 3(b) shows that the initial-state opacity fluctuates with increasing $\tau$ but does not reach the optimal initial-state opacity (0.329) achieved by our method.
These comparison experiments indicate that our proposed methods could exploit the noise in the observation to optimize opacity. This feature is not achievable using a policy computed from entropy-regularized MDPs, which do not employ the observation function.
6 Conclusion and Future Work
In this paper, we introduce information-theoretic definitions of opacity in dynamical systems modeled as Markov decision processes, against observers with imperfect information. Given the constraints on task performance, we develop opacity-enforcement policy gradient methods based on message passing in Hidden Markov Models (HMMs) for two basic types of opacity: last-state opacity and initial-state opacity. We expect the fundamental results on last-state opacity and initial-state opacity would be extended to other types of opacity, including language-based and model-based opacity and other obfuscation methods.
While the algorithm exhibits good performance in experiments, its drawback lies in high computational complexity due to the utilization of the forward-backward algorithm for gradient calculation. Finding alternative approaches to mitigate computation complexity will be among our future steps. Another direction to explore would be to investigate opacity-enforcement under different assumptions for the observer, for example, observers with imprecise knowledge about the agent’s policy. Lastly, though the objective is to maximize opacity, the same approaches can be used for minimizing opacity, and thus maximizing transparency, which is a desired property for AI and robotic systems interacting with humans. Future work would also look into potential applications and evaluate the effects of transparency in multi-agent teaming.
Acknowledgements
This work was sponsored in part by Army Research Laboratory under Cooperative Agreement Number W911NF-22-2-0233, the Army Research Office and was accomplished under Grant Number W911NF-22-1-0166 and in part by NSF under grant No. 2144113. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office, the Army Research Lab, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.